OpenStudy (anonymous):

Medal for the correct answer! Which of the following equations have the same solution? Explain each step in solving each equation using complete sentences. I. 4x – 14 = 2x + 8 II. 3x – 21 = 12 lll. x-2/5=4x+2/25

OpenStudy (anonymous):

I'll do the first one you find the rest! :D 4x – 14 = 2x + 8 ALL WE NEED TO DO IS FIND SOLUTION 4x - 2x + 14 + 8 2x = 22 x = 11 Understood? :) Now find the solution of the other equations. :) Brush up your Simultaneous equations skills a bit by going through KA videos. :)

OpenStudy (anonymous):

and the second one is 11 as well, just solve the equations and it will be easy. 3x = 33 x = 33/3

OpenStudy (anonymous):

lll. x-2/5=4x+2/25 please help with this one its my last one

OpenStudy (anonymous):

You can do it don't be lazy. if you have x -(2/5) = 4x + (2/25) you have to put every x on one side and the rest on the other. so you'll get 3x = -(2/5) - (2/25) now you do x = -0.16

OpenStudy (anonymous):

am i just pluging in for x now ?